Operations Report – February 2nd

Crew 221 Operations Report 2-2-2022 UPDATE 3

SOL: 0

Name of person filing report: Luis D. Monge

Non-nominal systems: Primary Generator, Hab lower deck 1 fluorescent light ballast (MDRS team in process of upgrade).

Notes on non-nominal systems:

Generator:

From what time last night: 21:00

To what time this morning: 07:00

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 79%

Diesel Reading – 25%

Station Propane Reading – 70%

Ethanol Free Gasoline – gallons

Water (loft tank): 50 gallons

Water Meter: units-

Water (static tank) – (MDRS team full refilling in progress)

Static to Loft Pump used – yes

Water in GreenHab –88.44 (MDRS team full refilling in progress)

Water in ScienceDome: 0 gallons

Toilet tank emptied: yes

Deimos rover used: no

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) 128.7

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 85%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: Yes

Hours: 134,8

Beginning charge: 100%

Ending charge: 80%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use:

Oil Added?

ATV Fuel Used: Gals

# Hours the ATVs were Used today:

Notes on ATVs:

HabCar used and why, where?

CrewCar used and why, where?

General Notes and Comments: full Cleaning and maintenance performed

Summary of the internet: Nominal.

Summary of Suit and Radios: Suits Nominal

Summary of Hab: Main fluorescent lights in Hab lower deck are being upgraded by MDRS team. New cameras installed and connected to existing station TV.

Summary of Science Dome operations: Nominal.

Summary of Ram operations: Nominal.

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: MDRS crew all healthy

This is the first ops report for Crew 221. Looking forward to a great time here!

General Notes and Comments: Cleaning and maintenance performed

Summary of the internet: Nominal.

Summary of Suit and Radios: 5 radios malfunctioning

Suits Nominal

Summary of Hab: (UPDATE) Main fluorescent lights in Hab lower deck are being upgraded by MDRS team. New stove installed. New cameras installed and connected to existing station TV. Water heater repaired and functioning.

Summary of Science Dome operations: Nominal.

Summary of Ram operations: Nominal. MDRS team began work to upgrade exterior outlets

Summary of any observatory issues: Nominal

Summary of Health and Safety issues: MDRS crew all healthy

This is the last ops report for Crew 220. Thanks for everything!